Find the p-value for the hypothesis test. A random sample of size 54 is taken. The sample has a mean of 426 and a standard deviation of 82.

H0: µ = 400

Ha: µ > 400

The p-value for the hypothesis test is .

Your answer should be rounded to 4 decimal places.
Answer #1

solution

This is the right tailed test .

The null and alternative hypothesis is ,

H0 :  \mu = 400

Ha : \mu > 400

Test statistic = z

= (\bar x - \mu ) / \sigma / \sqrt n

= (426-400) /52 / \sqrt 54

= 2.33

P(z >2.33 ) = 1 - P(z < 2.33) = 1 - 0.9901=0.0099

P-value = 0.0099